# Randomized Algorithms for Low-Rank Matrix and Tensor Decompositions

### v1

@@ 5.3 Tensor Decompositions
This work considers two fundamental formats of tensor decompositions: the canonical polyadic (CP) decomposition and the Tucker decomposition. We introduce them in Sect. 5.3 and 5.3, along with prototypical algorithms to compute them.
The CP decomposition has been known by many different names since its introduction in 1927: the polyadic form Hitchcock1927TheEO, PARAFAC (parallel factors) Harshman1970, CANDECOMP (canonical decomposition) carroll1970analysis; see also (KoldaBader2009, Table 3.1). It is often viewed as a higher-order extension of the matrix SVD KoldaBader2009.
The Tucker decomposition Tuck1963a; tucker1964; Tucker1966 is a representation of a tensor as a Multi-TTM product of a core tensor with a factor matrix for each mode. Like the CP decomposition, the Tucker decomposition has been known by several different names in the literature (KoldaBader2009, Table 4.1), and the Tucker decomposition can also be viewed as an extension of the matrix SVD, in the form of higher-order principal component analysis (PCA) kapteyn1986approach; kroonenberg1980principal; LeiboviciSabatier1998; Vasilescu2002.
We briefly summarize a few key properties of these decompositions in Sect. 5.3 before discussing low-rank tensor approximations that can be expressed in each format in Sect. 5.3.
@@
$\bm{\mathsf{X}}_{(j)}=\bm{\mathsf{A}}^{(j)}\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{\top}.$
@@
In practice, the CP decomposition is often computed by the alternating least squares (ALS) method carroll1970analysis; Harshman1970. Pseudocode for the general ALS procedure is given in Algorithm 5, which we quickly summarize.
Given a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, the ALS procedure outputs a CP decomposition $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\lambda};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$ with $R$ rank-one components, once some convergence criteria is met. This is accomplished through a sequence of linear least squares problems, where all but one factor matrix is fixed:
$\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\bm{\mathsf{X}}_{(j)}-\widehat{\bm{\mathsf{A}}}\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{\top}\right\|^{2}_{\text{F}}.$ (46)
@@
$\bm{\mathsf{A}}^{(j)}:=\bm{\mathsf{X}}_{(j)}\left[\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{\top}\right]^{\dagger},$ (47)
@@
If Algorithm 5 is naively implemented, the per-iteration cost to compute a CP decomposition is dominated by $N=\textup{size}(\bm{\mathcal{X}})$, which is computationally undesirable. Fortunately, since its introduction in 1970, CP-ALS has benefited from many modifications for improved efficiency, cf. Sect. 6.1 for recent algorithmic developments.
One strategy that has garnered attention for efficient CP-ALS incorporates a preprocessing step to compress the input tensor before executing Algorithm 5; this method is known as CANDELINC BroAndersson1998PartII; Carroll1980. The compression is often achieved through a Tucker decomposition, which we discuss next.
@@
6:   $\bm{\mathsf{A}}^{(j)}=\bm{\mathsf{X}}_{(j)}\left[\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{\top}\right]^{\dagger}$
@@
The Tucker decomposition expresses a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$ as a $d$-way Multi-TTM product:
@@
$\bm{\mathsf{X}}_{(j)}=\bm{\mathsf{A}}^{(j)}\bm{\mathsf{G}}_{(j)}\left(\bm{\mathsf{A}}^{(d)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(j+1)}\otimes\bm{\mathsf{A}}^{(j-1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(1)}\right)^{\top}.$ (49)
@@
A prototypical algorithm to compute the Tucker decomposition is Tucker’s “Method 1” Tucker1966, which has come to be known as the higher-order SVD (HOSVD) DeLathauwer2000a, summarized in Algorithm 6. For an input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots N_{d}}$, each factor matrix $\bm{\mathsf{A}}^{(j)}$ is computed as the leading-$R_{j}$ left singular vectors of the mode-$j$ unfolding $\bm{\mathsf{X}}_{(j)}$. If $R_{j}<\textup{rank}(\bm{\mathsf{X}}_{(j)})$, the decomposition is said to be “truncated.” When $R_{j}\ll N_{j}$, the Tucker decomposition is said to be a compressed representation of the input tensor. Once the factor matrices are computed, the core tensor is obtained as a Multi-TTM product of the original tensor with the factor matrix transposes. The computational cost of the HOSVD is $O(dn^{d+1}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$, letting $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots=R_{d}$ for notational simplicity.
For improved efficiency, the sequentially truncated HOSVD (ST-HOSVD) is often used Hackbusch2019; Vannieuwenhoven2012. In the ST-HOSVD, once a factor matrix is computed for mode $j$, the input tensor is compressed via mode-$j$ TTM before the next factor matrix is computed; i.e., for input tensor $\bm{\mathcal{X}}$, we first compute $\bm{\mathsf{A}}^{(1)}$ and form $\bm{\mathcal{Y}}=\bm{\mathcal{X}}\times_{1}\bm{\mathsf{A}}^{(1)}$, then compute $\bm{\mathsf{A}}^{(2)}$ from $\bm{\mathcal{Y}}$, then form $\bm{\mathcal{Z}}=\bm{\mathcal{Y}}\times_{2}\bm{\mathsf{A}}^{(2)}$, and so on. ST-HOSVD successively reduces the size of the tensor used in subsequent computations, and the core tensor is available immediately after computing the factor matrix for the last mode. The computational cost of ST-HOSVD is $O(\sum_{j=1}^{d}k^{j-1}n^{d-j+2}+k^{j}n^{d-j})$.
@@
Unlike the matrix SVD, the (ST-)HOSVD does not yield optimal approximation error. It has been shown that the relative error with (ST-)HOSVD is within $\sqrt{d}$ of an optimal Tucker decomposition of a given rank for a $d$-mode tensor, cf. (Ballard_Kolda_2025, Chapter 7). However, we can use the HOSVD output as a starting point for an ALS-style algorithm to compute Tucker decompositions, known as higher-order orthogonal iteration (HOOI) kroonenberg1980principal; DeLathauwer2000a. Pseudocode for HOOI is given in Algorithm 7.
As in CP-ALS, in HOOI, we seek to minimize the approximation error $\|\bm{\mathcal{X}}-\widehat{\bm{\mathcal{X}}}\|$, now over $\bm{\mathcal{G}}$ and $\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}$, subject to the constraints that $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1},\ldots,R_{d}}$ and that $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R_{j}}$ is orthogonal for $j=1,\ldots,d$. This tensorial minimization problem can be transformed into matrix least squares problems in each mode $j$, where we solve for each factor matrix via
$\bm{\mathsf{A}}^{(j)}=\arg\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R_{j}}}\left\|\left(\bm{\mathsf{A}}^{(d)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(j+1)}\otimes\bm{\mathsf{A}}^{(j-1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(1)}\right)\bm{\mathsf{G}}^{\top}_{(j)}\widehat{\bm{\mathsf{A}}}^{\top}-\bm{\mathsf{X}}^{\top}_{(j)}\right\|^{2}_{F}.$
Note the optimal solution to this least squares problem is given by the leading left singular vectors of each unfolding $\bm{\mathsf{X}}_{(j)}$ Andersson1998; DeLathauwer2000; Kolda2006. The core tensor can be computed as a column vector $\bm{\mathsf{g}}=\arg\min_{\bm{\mathsf{z}}\in\mathbb{F}^{\prod_{j}R_{j}\times 1}}\left\|\left(\bigotimes_{j=d}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{z}}-\bm{\mathsf{x}}\right\|^{2}_{2}$, where $\bm{\mathsf{x}}$ is the vectorization or unfolding of tensor $\bm{\mathcal{X}}$ into a large column vector of size $N\times 1$. To get $\bm{\mathcal{G}}$, we “re-fold” $\bm{\mathsf{g}}\in\mathbb{F}^{\prod_{j}R_{j}}$ back into a tensor of dimension $\mathbb{F}^{R_{1}\times\cdots\times R_{d}}$. If the factor matrices are computed as the leading $R_{j}$ left singular vectors of unfoldings, then the optimal core tensor is given by $\bm{\mathcal{G}}=\bm{\mathcal{X}}\times_{1}\left(\bm{\mathsf{A}}^{(1)}\right)^{\top}\times\cdots\times_{d}\left(\bm{\mathsf{A}}^{(d)}\right)^{\top}$. The per-iteration cost of Algorithm 7 is $O(dkn^{d}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$, letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$.
@@
6:   $\bm{\mathsf{A}}^{(j)}=\arg\min_{\bm{\mathsf{A}}\in\mathbb{F}^{N_{j}\times R_{j}}}\left\|\left(\bigotimes_{i=d,i\neq j}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{G}}^{\top}_{(j)}\bm{\mathsf{A}}^{\top}-\bm{\mathsf{X}}^{\top}_{(j)}\right\|^{2}_{F}$
@@
Mode ordering and TTM products with nonsingular matrices do not affect tensor rank. However, it is NP-hard to determine Hrastad1990.
@@
In practice, the target ranks for low-rank tensor decompositions are chosen heuristically. For example, we might choose the smallest rank $R$ for a CP decomposition that significantly reduces the relative error, in comparison to rank R1R-1, cf. (Ballard_Kolda_2025, Section 9.4.1). Generally, the optimization procedure in CP-ALS is repeated until the error no longer decreases or a maximum number of iterations is reached.
@@
For Tucker decompositions, it is useful to consider the $j$-rank of a tensor, denoted by $\textup{rank}_{j}(\bm{\mathcal{X}})$ and defined as the column rank of its mode-$j$ unfolding $\bm{\mathsf{X}}_{(j)}$ kruskal1989rank; DeLathauwer2000a. If $R_{j}=\textup{rank}_{j}(\bm{\mathcal{X}})$ for $j=1,\ldots,d$, then we say that $\bm{\mathcal{X}}$ is a rank-$(R_{1},\ldots,R_{d})$ tensor (distinct from its rank, defined above). We can compute an exact rank-$(R_{1},\ldots,R_{d})$ Tucker decomposition for any tensor $\bm{\mathcal{X}}$ with $R_{j}=\textup{rank}_{j}(\bm{\mathcal{X}})$ for $j=1,\ldots,d.$ However, if $R_{j}<\textup{rank}_{j}(\bm{\mathcal{X}})$ for any $j$, the decomposition is inexact and more difficult to compute. In practice, we often prescribe a core $j$-rank for $j=1,\ldots,d$ that meets some error tolerance or compression ratio (Ballard_Kolda_2025, Section 4.2). The (ST-)HOSVD algorithm has the nice property that it computes a rank-$(R_{1},\ldots,R_{d})$ decomposition with an approximation error that is within $\sqrt{d}$ of the optimal rank-$(R_{1},\ldots,R_{d})$ approximation error.
@@
Whether it is better to represent an input tensor in the CP or Tucker format is generally problem-dependent. Under mild assumptions, the CP decomposition of a tensor is unique, up to scaling and permutation of its rank-one components tenberge2002uniqueness; harshman1984data; Harshman1970; kruskal1977three; kruskal1989rank; sidiropoulos2000uniqueness. The same cannot be said for Tucker decompositions; the mode-$j$ factor matrix can be multiplied by a nonsingular matrix and the core tensor updated via mode-$j$ product of the original core tensor with the inverse of the nonsingular matrix. As a result, the factors of a Tucker decomposition are less interpretable in the context of the original data than those of a CP decomposition, cf. (Ballard_Kolda_2025, Section 9.6–9.8).
However, as previously mentioned, the Tucker decomposition is highly advantageous for data compression, even as a preprocessing step for a CP decomposition, e.g., CANDELINC BroAndersson1998PartII; Carroll1980. Despite its utility in data compression, the Tucker decomposition still suffers from the curse of dimensionality: letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$, the cost of storing a Tucker decomposition is $O(k^{d}+dkn)$. By contrast, the storage cost of a CP decomposition is linear in its rank, but its rank may be very large to achieve the desired approximation accuracy.
@@
In short, the notion of rank for tensor decompositions is more nuanced than for matrices, and the choice of representation and target rank (or $j$-rank) should be informed by the desired trade-off between approximation accuracy and computational performance. For more detailed treatments, we refer the reader to Ballard_Kolda_2025; DeLathauwer2000a; Grasedyck2013; KoldaBader2009.
@@
In the previous sections, we introduced two fundamental tensor representations, the CP and the Tucker decomposition, and we summarized their key properties. In this section, we survey low-rank tensor factorizations that can be expressed in each format, which we characterize primarily by whether or not the factor matrices are orthogonal. Namely, we assume that the standard Tucker decomposition has orthonormal factor matrices, whereas CP decompositions generally do not. Higher-order versions of ID/CUR may be expressed using orthogonal factor matrices or factor matrices that contain entries of the input tensor, so we further characterize them by whether the core subtensor or the factor matrices contain entries of the original tensor.
@@
As in the matrix setting (Sect. 2.4), a tensor ID or CUR factorization comprises actual entries of the input tensor. The advantages of ID/CUR factorizations for tensors are similar to those for matrices. The tensor ID/CUR factorization can be interpreted more easily in the context of the original tensorial data than factorizations involving multi-linear transformations of the data. Moreover, the ID/CUR preserves the structure of the input tensor (e.g., sparsity or non-negativity), and it can be computed efficiently with minimal storage requirements.
@@
The work of MahoneyMaggioniDrineas2008_tensorCUR that introduced “tensor CUR factorizations” falls into the first category. In their algorithm, a $\textup{rank-}(N_{1},N_{2},R)$ core tensor is constructed by selecting $R$ indices from one mode and keeping all indices from the other two. Core tensor skeleton selection methods for 3-mode tensor CUR were generalized to $d$-mode tensors in Caiafa2010, for Tucker decompositions with the same $j$-rank in each mode. The “structure-preserving ST-HOSVD” of MinsterSaibabaKilmer2020, initially developed for sparse rank-$(R_{1},\ldots,R_{d})$ tensors, may also be interpreted as a tensor CUR factorization of the first category. The “Chidori” and “Fiber” tensor CUR factorizations with rank-$(R_{1},\ldots,R_{d})$ core tensors were characterized in Cai2021, inspired by those in Caiafa2010; these factorizations were utilized for robust tensorial principal component analysis in Cai2024. The recently proposed “CoreID” in Zhang2025 is similar to the CUR factorizations of Cai2021. In all of these methods, the entries of the associated factor matrices are computed based on the specific form of the factorization and on the skeletons selected for the core tensor, but the factor matrices are not themselves sub-matrices of unfoldings.
The second category above comprises those tensor ID/CUR factorizations in which the entries of the factor matrices are entries of the input tensor. The early algorithm of Drineas2007tensorSVD to compute a $\textup{rank-}(R_{1},\ldots,R_{d})$ “approximate tensor SVD” forms the factor matrices in a Tucker decomposition as randomly sampled columns of unfoldings; this factorization would later be termed a “higher-order ID” in Saibaba2016, where the mode-$j$ factor matrix in a Tucker decomposition is computed as column skeletons of the mode-$j$ unfolding selected by greedy pivoting. In FriedlandMehrmannMiedlarNkengla2011, an HOOI procedure is developed for 3-mode Tucker tensors that yields factor matrices of original tensor entries, and the “CURT” algorithm of ZhaoSong2018RETL; Song2019 computes tensor CUR decompositions in the Tucker format with strong accuracy guarantees. The CURT algorithm was recently generalized in Zhang2025 to compute rank-$(R_{1},\ldots,R_{d})$ Tucker decompositions where the core tensor is unconstrained and the factor matrices are column skeletons of mode unfoldings; this algorithm is also suitable for input tensors expressed in the CP format. The tensor ID algorithm of Biagioni2015 computes a CP decomposition of a $d$-mode tensor, in which factor matrices are comprised of entries of the input tensor, as does the more recent work of MalikBecker2020. We detail several algorithms that leverage randomization to compute tensor ID/CUR factorizations efficiently in Sect. 6.

### v2

@@ 5.3 Tensor Decompositions
This work considers two fundamental formats of tensor decompositions: the canonical polyadic (CP) decomposition and the Tucker decomposition. The CP decomposition has been known by many different names since its introduction in 1927 in Hitchcock1927TheEO; Hitchcock1927Multiple, cf. (KoldaBader2009, Table 3.1). This decomposition is often viewed as a higher-order extension of the matrix SVD, as it is formed from rank-one components. The Tucker decomposition Tuck1963a; tucker1964; Tucker1966 is a representation of a tensor as a Multi-TTM product of a core tensor with a factor matrix for each mode. Like the CP decomposition, the Tucker decomposition has been known by several different names in the literature, cf. (KoldaBader2009, Table 4.1), and the Tucker decomposition can also be viewed as an extension of the matrix SVD, in the form of higher-order principal component analysis (PCA).
@@
$\bm{\mathsf{X}}_{(j)}=\bm{\mathsf{A}}^{(j)}\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{*}.$
@@
In practice, the CP decomposition is often computed by the alternating least squares (ALS) method carroll1970analysis; Harshman1970 in Algorithm 5. Given a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, CP-ALS outputs $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\lambda};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$ with $R$ rank-one components through a sequence of linear least squares problems, where all but one factor matrix is fixed:
$\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\bm{\mathsf{X}}_{(j)}-\widehat{\bm{\mathsf{A}}}\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{*}\right\|^{2}_{\text{F}}.$ (40)
@@
$\bm{\mathsf{A}}^{(j)}:=\bm{\mathsf{X}}_{(j)}\left[\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{*}\right]^{\dagger},$ (41)
@@
If Algorithm 5 is naively implemented, the per-iteration cost to compute a CP decomposition is dominated by $N=\textup{size}(\bm{\mathcal{X}})$, which is computationally undesirable. Fortunately, since its introduction in 1970, CP-ALS has benefited from many modifications for improved efficiency. One strategy that has garnered attention incorporates a preprocessing step to compress the input tensor before executing Algorithm 5, known as CANDELINC BroAndersson1998PartII; Carroll1980, which may be achieved through a Tucker decomposition.
@@
6:   $\bm{\mathsf{A}}^{(j)}=\bm{\mathsf{X}}_{(j)}\left[\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{*}\right]^{\dagger}$
@@
The Tucker decomposition expresses a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$ as a $d$-way Multi-TTM
@@
$\bm{\mathsf{X}}_{(j)}=\bm{\mathsf{A}}^{(j)}\bm{\mathsf{G}}_{(j)}\left(\bm{\mathsf{A}}^{(d)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(j+1)}\otimes\bm{\mathsf{A}}^{(j-1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(1)}\right)^{*}.$ (43)
@@
A prototypical algorithm to compute the Tucker decomposition is Tucker’s “Method 1” Tucker1966, which has come to be known as the higher-order SVD (HOSVD) DeLathauwer2000a, summarized in Algorithm 6. For an input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots N_{d}}$, each factor matrix $\bm{\mathsf{A}}^{(j)}$ is computed as the leading-$R_{j}$ left singular vectors of the mode-$j$ unfolding $\bm{\mathsf{X}}_{(j)}$. If $R_{j}<\textup{rank}(\bm{\mathsf{X}}_{(j)})$, the decomposition is said to be “truncated.” When $R_{j}\ll N_{j}$, we say the tensor is compressed. The core tensor is a Multi-TTM of the original tensor with the factor matrix transposes. The total cost is $O(dn^{d+1}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$, letting $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots=R_{d}$ for notational simplicity.
For improved efficiency, the sequentially truncated HOSVD (ST-HOSVD) is often used Vannieuwenhoven2012. In the ST-HOSVD, once the mode-$j$ factor matrix is computed, the tensor is compressed via mode-$j$ TTM before the next factor matrix is computed; i.e., we first compute $\bm{\mathsf{A}}^{(1)}$ and form $\bm{\mathcal{Y}}=\bm{\mathcal{X}}\times_{1}\bm{\mathsf{A}}^{(1)}$, then compute $\bm{\mathsf{A}}^{(2)}$ from $\bm{\mathcal{Y}}$, then form $\bm{\mathcal{Z}}=\bm{\mathcal{Y}}\times_{2}\bm{\mathsf{A}}^{(2)}$, and so on. ST-HOSVD successively reduces the size of the tensor used in subsequent computations, and the core tensor is available immediately after computing the factor matrix for the last mode. The total cost is $O(\sum_{j=1}^{d}k^{j-1}n^{d-j+2}+k^{j}n^{d-j})$.
@@
Unlike the matrix SVD, the (ST-)HOSVD does not yield optimal approximation error. It has been shown that the relative error with (ST-)HOSVD is within $\sqrt{d}$ of an optimal Tucker decomposition of a given rank for a $d$-mode tensor, cf. (Ballard_Kolda_2025, Chapter 7). However, we can use the HOSVD output as a starting point for an ALS-style algorithm to compute Tucker decompositions, known as higher-order orthogonal iteration (HOOI) or Tucker-ALS kroonenberg1980principal; DeLathauwer2000a. Pseudocode for HOOI is given in Algorithm 7. We seek to minimize the approximation error $\|\bm{\mathcal{X}}-\widehat{\bm{\mathcal{X}}}\|$, now over $\bm{\mathcal{G}}$ and $\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}$, subject to the constraints that $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1},\ldots,R_{d}}$ and that $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R_{j}}$ is orthogonal for $j=1,\ldots,d$. This minimization problem can be transformed into a matrix least squares problem in each mode $j$, so that
$\bm{\mathsf{A}}^{(j)}=\arg\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R_{j}}}\left\|\left(\bm{\mathsf{A}}^{(d)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(j+1)}\otimes\bm{\mathsf{A}}^{(j-1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(1)}\right)\bm{\mathsf{G}}^{*}_{(j)}\widehat{\bm{\mathsf{A}}}^{*}-\bm{\mathsf{X}}^{*}_{(j)}\right\|^{2}_{F}.$
The solution to this least squares problem is given by the leading left singular vectors of each unfolding $\bm{\mathsf{X}}_{(j)}$ DeLathauwer2000; Kolda2006. The core tensor can be computed as a column vector $\bm{\mathsf{g}}=\arg\min_{\bm{\mathsf{z}}\in\mathbb{F}^{\prod_{j}R_{j}\times 1}}\left\|\left(\bigotimes_{j=d}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{z}}-\bm{\mathsf{x}}\right\|^{2}_{2}$, where $\bm{\mathsf{x}}$ is the vectorization or unfolding of tensor $\bm{\mathcal{X}}$ into a large column vector of size $N\times 1$. To get $\bm{\mathcal{G}}$, we “re-fold” $\bm{\mathsf{g}}\in\mathbb{F}^{\prod_{j}R_{j}}$ back into a tensor of dimension $\mathbb{F}^{R_{1}\times\cdots\times R_{d}}$. If the factor matrices are computed as the leading $R_{j}$ left singular vectors of unfoldings, then the optimal core tensor is given by $\bm{\mathcal{G}}=\bm{\mathcal{X}}\times_{1}\left(\bm{\mathsf{A}}^{(1)}\right)^{*}\times\cdots\times_{d}\left(\bm{\mathsf{A}}^{(d)}\right)^{*}$. The per-iteration cost of Algorithm 7 is $O(dkn^{d}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$, letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$.
@@
6:   $\bm{\mathsf{A}}^{(j)}=\arg\min_{\bm{\mathsf{A}}\in\mathbb{F}^{N_{j}\times R_{j}}}\left\|\left(\bigotimes_{i=d,i\neq j}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{G}}^{*}_{(j)}\bm{\mathsf{A}}^{*}-\bm{\mathsf{X}}^{*}_{(j)}\right\|^{2}_{F}$
@@
Mode ordering and TTM products with nonsingular matrices do not affect tensor rank. However, it is NP-hard to determine Hrastad1990. In practice, the target ranks for low-rank tensor decompositions are chosen heuristically. For example, we might choose the smallest rank $R$ for a CP decomposition that significantly reduces the relative error, in comparison to rank R1R-1, cf. (Ballard_Kolda_2025, Section 9.4.1). Generally, the optimization procedure in CP-ALS is repeated until the error no longer decreases significantly.
@@
For Tucker decompositions, it is useful to consider the $j$-rank of a tensor, denoted by $\textup{rank}_{j}(\bm{\mathcal{X}})$ and defined as the column rank of its mode-$j$ unfolding $\bm{\mathsf{X}}_{(j)}$ kruskal1989rank; DeLathauwer2000a. If $R_{j}=\textup{rank}_{j}(\bm{\mathcal{X}})$ for $j=1,\ldots,d$, then we say that $\bm{\mathcal{X}}$ is a rank-$(R_{1},\ldots,R_{d})$ tensor (distinct from its rank, defined above). We can compute an exact rank-$(R_{1},\ldots,R_{d})$ Tucker decomposition for any tensor $\bm{\mathcal{X}}$ with $R_{j}=\textup{rank}_{j}(\bm{\mathcal{X}})$ for $j=1,\ldots,d.$ However, if $R_{j}<\textup{rank}_{j}(\bm{\mathcal{X}})$ for any $j$, the decomposition is inexact and more difficult to compute. In practice, we often prescribe a core $j$-rank for $j=1,\ldots,d$ that meets some error tolerance or compression ratio (Ballard_Kolda_2025, Section 4.2).
@@
Whether it is better to represent an input tensor in the CP or Tucker format is generally problem-dependent. Under mild assumptions, the CP decomposition of a tensor is unique, up to scaling and permutation of its rank-one components tenberge2002uniqueness; harshman1984data; Harshman1970; kruskal1977three; kruskal1989rank; sidiropoulos2000uniqueness. The same cannot be said for Tucker decompositions; the mode-$j$ factor matrix can be multiplied by a nonsingular matrix and the core tensor updated via mode-$j$ TTM with its inverse. As a result, the factors of a Tucker decomposition are less interpretable in the context of the original data than those of a CP decomposition.
@@
Despite its utility in data compression, the Tucker decomposition still suffers from the curse of dimensionality: letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$, the cost of storing a Tucker decomposition is $O(k^{d}+dkn)$. By contrast, the storage cost of a CP decomposition is linear in its rank, but its rank may be very large to achieve the desired accuracy. In short, the notion of rank for tensor decompositions is more nuanced than for matrices, and the choice of representation and target rank (or $j$-rank) should be informed by the desired trade-off between approximation accuracy and computational performance. For more detailed treatments, we refer the reader to Ballard_Kolda_2025; DeLathauwer2000a; Grasedyck2013; KoldaBader2009.
@@
In the previous sections, we introduced two fundamental tensor representations, the CP and the Tucker decomposition, and we summarized their key properties. In this section, we survey low-rank tensor factorizations that can be expressed in each format.
@@
As in the matrix setting (Sect. 2.5), a tensor ID or CUR factorization comprises actual entries of the input tensor. The advantages of ID/CUR factorizations for tensors are similar to those for matrices. The tensor ID/CUR factorization can be interpreted more easily in the context of the original tensorial data than factorizations involving multi-linear transformations of the data. Moreover, the ID/CUR preserves the structure of the input tensor (e.g., sparsity or non-negativity), and it can be computed efficiently with minimal storage requirements.
@@
The work of MahoneyMaggioniDrineas2008_tensorCUR that introduced “tensor CUR factorizations” falls into the first category. In their algorithm, a $\textup{rank-}(N_{1},N_{2},R)$ core tensor is constructed by selecting $R$ indices from one mode and keeping all indices from the other two. Core tensor skeleton selection methods for 3-mode tensor CUR were generalized to $d$-mode tensors in Caiafa2010, for Tucker decompositions with the same $j$-rank in each mode. The “structure-preserving ST-HOSVD” of MinsterSaibabaKilmer2020, initially developed for sparse rank-$(R_{1},\ldots,R_{d})$ tensors, may also be interpreted as a tensor CUR factorization of the first category. The “Chidori” and “Fiber” tensor CUR factorizations with rank-$(R_{1},\ldots,R_{d})$ core tensors were characterized in Cai2021, inspired by those in Caiafa2010. The recently proposed “CoreID” in Zhang2025 is similar to the CUR factorizations of Cai2021. In all of these methods, the entries of the associated factor matrices are computed based on the specific form of the factorization and on the skeletons selected for the core tensor, but the factor matrices are not themselves sub-matrices of unfoldings.
The second category above comprises those tensor ID/CUR factorizations in which the entries of the factor matrices are entries of the input tensor. The early algorithm of Drineas2007tensorSVD to compute a $\textup{rank-}(R_{1},\ldots,R_{d})$ “approximate tensor SVD” forms the factor matrices in a Tucker decomposition as randomly sampled columns of unfoldings; this factorization would later be termed a “higher-order ID” in Saibaba2016, where the mode-$j$ factor matrix in a Tucker decomposition is computed as column skeletons of the mode-$j$ unfolding selected by greedy pivoting. We detail several recent algorithms that leverage randomization to compute tensor ID/CUR approximations efficiently in Sect. 6.